# Search-Based Software Engineering in the Landscape of AI Foundation Models

### v1

@@ 6 Related Works
To the best of our knowledge, this paper presents the first forward-looking roadmap for SBSE in the context of FMs. While several existing studies have presented the state of research and future directions, many are outdated or no longer reflect the latest advancements, specifically the ones due to FMs. Our work complements the existing works by introducing a timely and forward-thinking vision for the evolution of SBSE in the FM era. Initially, Harman et al. [19] presented a detailed survey of the SBSE research state till 2010, challenges, and future directions. Subsequently, Colanzi et al. [11] presented a systematic mapping study to analyze the evolution of SBSE from 2009 to 2019. Sarro [34] later presented a brief vision highlighting SBSE applications for building responsible software systems. Recently, Wu et al. [42] presented a literature review and a roadmap for integrating LLMs with evolutionary algorithms (EAs). In comparison, our work focuses on the synergistic potential between SBSE, which leverages EAs for solving SE problems, and FMs, which encompass not only LLMs but also VLMs, SAMs, and MMs. Furthermore, we specifically present a roadmap focusing on the potential interplay between SBSE and FMs within the SE context, whereas their work focused on cross-domain research areas.
@@
In the context of SBST, McMinn [26] conducted a comprehensive survey that focused on test data generation using search techniques. Similarly, Ali et al. [2] provided a systematic review of the literature on empirical studies related to search-based test case generation up to the year 2007. McMinn [27] presented a review of the SBST area, including studies till 2011, open challenges, and future directions. Neelofar et al. [29] analyzed the strengths and limitations of automated techniques in SBST, specifically focusing on instance space analysis. In recent work, Fraser and Arcuri [14] outlined a brief vision for the future of search-based test generation in the context of LLMs. In contrast, our work takes a broader perspective, exploring the interplay between SBSE and FMs (in addition to LLMs).
7 Conclusion and Future Work
@@
Our analysis of the current state of the art revealed that most research focuses on applying large language models (LLMs) in evolutionary computation and search-based software engineering (SBSE), leaving many research opportunities to explore other foundation models (FMs) like vision language models and multimodal FMs. We also noted some initial, yet immature, efforts in using LLMs for SBSE, SBSE for LLMs, and their combined application to solve various software engineering problems. Building on these insights, we presented a roadmap outlining forward-looking opportunities and research directions, focusing on the potential synergy between FMs and SBSE. In the future, we plan to extend this roadmap in two directions: (i) enhancing the potential of SBSE and FMs across the entire software engineering lifecycle and (ii) exploring application-specific synergies between SBSE and FMs in emerging domains such as quantum computing, extended reality, and autonomous systems.
@@ Acknowledgments
This research is supported by the RoboSAPIENS project, funded through the European Commission Horizon Europe program under Grant Agreement No. 101133807.

### v2

@@ 11 Related Work
To the best of our knowledge, this paper presents the first forward-looking roadmap for SBSE in the context of FMs. While several existing studies have presented the state of research and future directions, many are outdated or no longer reflect the latest advancements, particularly those related to FMs. Our work complements existing work by introducing a timely and forward-thinking vision for the evolution of SBSE in the FM era. Initially, Harman et al. [68] presented a detailed survey of the SBSE research state till 2010, challenges, and future directions. Subsequently, Colanzi et al. [33] presented a systematic mapping study to analyze the evolution of SBSE from 2009 to 2019. Sarro [147] later presented a brief vision highlighting SBSE applications for building responsible software systems. Recently, Wu et al. [193] presented a literature review and a roadmap for integrating LLMs with evolutionary algorithms (EAs). In comparison, our work focuses on the synergistic potential between SBSE, which leverages EAs for solving SE problems, and FMs, which encompass not only LLMs but FMs more broadly. We specifically present a roadmap that explores the interplay between SBSE and FMs within the context of SE, distinguishing it from their work, which focused on cross-domain research areas. In addition, we discuss the challenges of empirical evaluations, analyze the disruptive effect of FMs on SBSE, and identify research opportunities in emerging domains.
@@
In the context of SBST, McMinn [112] conducted a comprehensive survey that focused on test data generation using search techniques. Similarly, Ali et al. [2] provided a systematic review of the literature on empirical studies related to search-based test case generation up to the year 2007. McMinn [113] presented a review of the SBST area, including studies till 2011, open challenges, and future directions. Later, Harman et al. [69] reviewed the SBST literature up to 2014, analyzed the growth of the field, and highlighted open challenges. In another study, Neelofar et al. [121] analyzed the strengths and limitations of automated techniques in SBST, specifically focusing on instance space analysis. More recently, Fraser and Arcuri [53] outlined a brief vision for the future of search-based test generation in the context of LLMs. In contrast, our work takes a broader perspective, exploring the interplay between SBSE and FMs (beyond just LLMs) from different aspects, challenges of empirical evaluations involving their integrated use, the disruptive effect of FMs on SBSE, and research opportunities in emerging domains.
There are some literature reviews analyzing domain-specific applications of SBSE. Ramirez et al. [137] presented a systematic review focusing on available approaches to make an interactive search process by adding humans in the loop. Mazzonetto et al. [109] conducted a systematic mapping study to analyze the state of research targeting SBSE for enterprise applications. Khoshnevis and Ardestani [82] presented a systematic review of SBSE applications in the software product line engineering domain. Compared to these works, we propose a forward-looking roadmap that highlights open challenges and research opportunities, focusing on the synergies between SBSE and FMs, their integration challenges, the potential of FMs to transform SBSE, and advancing research in emerging domains.
12 Conclusion
@@
Our analysis of the current state of the art revealed that most research focuses on applying large language models (LLMs) in evolutionary computation and search-based software engineering (SBSE), leaving many research opportunities to explore other FMs, like vision language models and multimodal FMs. We also noted some initial, yet immature, efforts in using LLMs for SBSE, SBSE for LLMs, and their combined application to solve various software engineering problems. Building on these insights, we presented a roadmap outlining forward-looking opportunities and research directions, focusing on the potential synergy between SBSE and FMs. In this context, we analyzed the current research landscape, existing challenges, and promising research opportunities, organizing our analysis around five key aspects: (1) leveraging FMs for SBSE design, (2) applying FMs to complement SBSE in SE problems, (3) utilizing SBSE to address FM challenges, (4) adapting SBSE practices for FMs tailored to SE activities, and (5) exploring the interplay between SBSE and FMs. For empirical evaluations involving the integrated use of SBSE and FMs, we presented key challenges that require further attention from the research community. Furthermore, we explored the future of SBSE in the context of FMs using McLuhan’s Tetrad and identified research opportunities to take advantage of their synergistic potential for addressing challenges in emerging domains, covering autonomous driving systems, robotics, AI-enabled systems, Internet of Things, and quantum computing.
@@ Acknowledgments
This research is supported by the RoboSAPIENS project, funded through the European Commission’s Horizon Europe program under Grant Agreement No. 101133807. Andrea Arcuri is funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (EAST project, grant agreement No. 864972). Paolo Arcaini is supported by the ASPIRE grant No. JPMJAP2301, JST.